# Mitigating Polarization in Recommender Systems via Network-aware Feedback Optimization

### v1

@@ 3 Recommender system design
We approach Problem 1 by designing a dynamic feedback controller inspired by the projected-gradient descent algorithm in [38, Section 3A]. The resulting recommender system dynamically generates recommendations as
@@
    Real-time opinions and users’ clicking behaviour estimation via supervised learning;
@@ 3.1 Level I: Opinions & Clicking Behaviour Estimation
The opinions and clicking behaviours are estimated using an Artificial Neural Network (ANN). Note that the use of the ANN allows us to take into account the dynamics of opinions thus circumventing the problem of getting access to online state measurements.
We estimate the opinions by relying on implicit feedback from the users, i.e. the CTR $y\approx\mathbb{E}\big{[}\mathcal{B}(g(p,x))\big{]}$ of the users on recommendations with position $p$. Thus, we state that there exists a map to extract the opinions from the clicks and the positions, $x=\beta(y,p)$. To bound the resulting opinions and clicking behaviour estimation error from the ANN, we work under the following regularity assumptions [17, 39].
@@
    the continuous map $\beta:[0,1]^{n}\times[-1,1]^{n}\to[-1,1]^{n}$ is such that $\beta(y,p)=x+\theta(x)$ with $\lVert\theta(x)\rVert\leq\theta_{x}$, for all $x\in\mathcal{X}\subseteq[-1,1]^{n}$, and $\theta_{x}<\infty$. Moreover, the image $\beta(\mathcal{Y},\mathcal{P})$ is compact for any $\mathcal{Y}\subseteq[0,1]^{n},\mathcal{P}\subseteq[-1,1]^{n}$;
@@
    there exists a constant $\alpha_{y}<\infty$ such that the composite function $g(p,\beta(y,p))=y+\nabla_{x}g(p,x)^{\top}\theta(x)+\alpha(y)$, with $\lVert\alpha(y)\rVert\leq\alpha_{y}$, for all $y\in\mathcal{Y}$. $\square$
Assumption 3 i) guarantees the map $\beta(y,p)$ to provide an estimate on users’ opinions upper bounded by $\theta_{x}$. Similarly, Assumption 3 ii), guarantees the (opinion) map $\beta(y,p)$ to (implicitly) provide an estimate on clicking behaviour upper bounded by $M_{x}\theta_{x}+\alpha_{y}$, with $M_{x}$ resulting from Assumption 2. We will qualitatively describe how these upper bound errors can be tightened in Section 3.4.
@@ 3.1.1 Opinion Estimation
For each user $i\in[n]$, the input and output training sets are $\begin{bmatrix}\mathcal{X}_{p}\\ \mathcal{X}_{y,i}\end{bmatrix}\in\begin{Bmatrix}[-1,1]^{n\times m}\\ [0,1]^{1\times m}\end{Bmatrix}$, $\mathcal{X}_{x,i}\in[-1,1]^{1\times m}$, respectively.
Given that users interact with each other, the recommendations to users connected to user $i$ will indirectly influence user $i$’s opinion $x_{i}$. However, $x_{i}$ is not directly affected by the CTR of other users. To model this, we design the input layer with $n+1$ neurons for each user: $n$ neurons associated with $p$ and one neuron associated with $y_{i}$. The output layer comprises a single neuron to represent $x_{i}$. Additionally, we incorporate an intermediate layer with $n+2$ neurons. The intermediate layer employs a hyperbolic tangent activation function $\tanh(x):\mathbb{R}\rightarrow[-1,1]$, while the output layer uses an identity activation function.
@@
Under Assumption 3, the opinion estimation error is upper bounded in the $\ell_{2}$-norm as $\lVert e_{x}\rVert\leq\epsilon_{x}$, with
@@
where $\underset{y\in\mathcal{X}_{y},p\in\mathcal{X}_{p}}{\mathrm{sup}}\lVert\beta(y,p)-\hat{\beta}(y,p)\rVert_{\infty}$ represents the maximum opinion estimation error during training of the ANN, $\omega_{\beta_{i}}(\gamma_{x})$ denotes the minimum modulus of continuity of $\beta_{i}$ on the training set as in Definition 1, and $v_{0,i}$ represents the bias from the hidden layer to the output layer of the ANN on user $i$.
@@ 3.1.2 Clicking Behaviour Estimation
For each user $i\in[n]$, the input and output training sets are $\begin{bmatrix}\mathcal{X}_{p,i}\\ \mathcal{X}_{x,i}\end{bmatrix}\in[-1,1]^{2\times m}$, $\mathcal{X}_{y,i}\in[0,1]^{1\times m}$, respectively. We then construct an ANN similarly as for opinion estimation [^1] We omit the details due to space limitations and refer the interested reader to for a detail discussion.. {arxiv} To estimate the clicking behaviour, we proceed similarly as for opinion estimation. For each user $i\in[n]$, the input and output training sets are $\begin{bmatrix}\mathcal{X}_{p,i}\\ \mathcal{X}_{x,i}\end{bmatrix}\in[-1,1]^{2\times m}$, $\mathcal{X}_{y,i}\in[0,1]^{1\times m}$, respectively. Given the recommender system’s personalization of positions to user’s interests, the $i$-th user’s CTR $y_{i}$ only depends on their steady-state opinion $x_{i}$ and the provided position $p_{i}$. Hence, we have two input neurons ($x_{i},p_{i}$) and one neuron in the output layer ($y_{i}$). Further, we use $3$ intermediate layers with $5$ neurons. The hyperbolic tangent and the identity are used as activation functions at the intermediate layers and the output layer, respectively.
@@
Under Assumptions 23, the clicking behaviour estimation error is upper-bounded in the $\ell_{2}$-norm as $\lVert e_{y}\rVert\leq\epsilon_{g}$, with
@@ 3.2 Level II: Online Sensitivity Learning
To estimate the sensitivity $\nabla_{p}h(p,d)$ in (7), we develop a Kalman filter, similar to [16]. Given that opinion dynamics and recommendations evolve over time scales that are comparable in magnitude, in order to make our OFO scheme [40] applicable, we introduce an auxiliary mechanism that guarantees time-scale separation between the opinion dynamics and the recommender system, see Remark 1.
@@ 3.2.1 Kalman Filter for Sensitivity Learning
The measurement model (11) is obtained from a first-order Taylor approximation of the steady-state map around $p^{k-1}$, i.e. $h(p^{k},d)\approx h(p^{k-1},d)+\nabla_{p}h(p^{k-1},d)^{\top}(p^{k}-p^{k-1})$. With this intuition, it can be seen that the measurement model is a linear time-variant approximation of the temporal evolution of the sensitivity $\nabla_{p}h(p^{k},d)$. It is to be noted that in reality, the steady-state $h(p^{k},d)$ cannot be perfectly achieved. Thus, it is common practice to use $\Delta x^{k+1,k}_{ss}=x^{k+1}-x^{k}$ as the measurement in (11), see [16].
@@
The trigger mechanism is introduced for the recommender system to take action at stationarity, see Remark 1, and to collect a sufficient number of clicks to ensure a proper estimate of the CTR needed for opinion estimation. The trigger mechanism is turned on at integer multiples $i\in\mathbb{N}$ of the time period $T$. We define $\mathcal{T}$ as the set of all trigger time instances, with $\tau_{i}\in\mathcal{T}$ representing the most recent trigger time instance prior to $k$. Note that the update (3.2.1) uses the opinion estimate $\hat{x}$ introduced in §3.1.1 rather than the real users’ opinion, with $\hat{x}^{k+1}=\hat{\beta}(y^{k},p^{k})$, where $y^{k}:=\sum_{t=\tau_{i}}^{k}c^{t}/(k-\tau_{i}+1)$ is the CTR since the most recent trigger time instant. Finally, the Kalman filter gain $K^{k}$ in (3.2.1) is given by
@@
The assumption of white noise for the process model is standard in the context of sensitivity learning in feedback optimization [16]. Intuitively, the process perturbation determines the degree of trust one puts on the sensitivity estimates. It is also reasonable to assume that the external influence is uncorrelated among users in certain cases, for example the extended FJ model in (2) where $d\equiv x_{0}$. We also note that since $\hat{x}$ is estimated using the ANN, the steady-state estimation error $e_{x}$ is not correlated with the process or measurement noises in (10), (11). Under Assumption 4, the covariances simplify as $Q^{k}=(\sigma_{q}^{k})^{2}I_{n^{2}}$ and $R^{k}=(\sigma_{r}^{k})^{2}I_{n}$, for some $\sigma_{q}$, $\sigma_{r}>0$. The tuning of $\sigma_{q},\sigma_{r}$, is based on heuristics and is described in Appendix 5.4.
@@ 3.3.1 Gradient Estimation
where $\nabla_{j}\hat{\varphi}^{\rm ctr}_{i},j\in\{p,x\}$ denotes the $i$-th entry of the gradient, $e_{i}\in\mathbb{R}^{n}$ refers to the $i$-th vector of the canonical basis of $\mathbb{R}^{n}$, and $\mu$ is a smoothing parameter. The cost $\hat{\varphi}^{\rm ctr}(p,x)=-1_{n}^{\top}\hat{g}(p,x)$ is evaluated as described in §3.1.2.
@@ 3.4 Closed-loop Convergence Guarantees
We provide convergence guarantees for the sensitivity estimation process (3.2.1) and for the closed-loop interconnection between the opinions and recommendations. Before stating the closed-loop convergence theorems, we make some important observations regarding opinion estimation.
With our problem setup, the measurement information at each time instant (i.e. click/ no click) is not rich enough to produce accurate real-time opinion estimation at every time instant. As shown in Algorithm 1, opinion estimation is carried out using the CTR recorded on recommendations with a constant position since the previous trigger instant. Hence, in light of Assumption 1, the resulting opinion estimate at the end of each trigger period is an estimate of the steady-state map $h(p^{k},d)$, thus differing from real-time state estimation [17]. By considering a longer trigger period $T$, the CTR at the end of a period turns into a better approximation of $\mathbb{E}\big{[}g(p,x)\big{]}$. We qualitatively state that the modelling errors $\theta_{x},\alpha_{y}$ in Lemma 2,3.2 reduces on increasing $T$. Furthermore, as the steady-state opinions $h(p^{k},d)$ are directly estimated at each trigger instant, we circumvent the need to separately prove exponential convergence for varying inputs on the opinion dynamics.
@@
To quantify performance on the recommendations, we use the fixed-point residual mapping [42, Eq. (5)]
$\mathcal{G}(p):=\frac{1}{\eta}\Big{(}p-\Pi_{[-1,1]^{n}}\big{[}p-\eta\Phi(p,h(p,d))\big{]}\Big{)},$ (18)
The fixed-point residual mapping is zero at a critical point of (5), and is a common metric to quantify convergence of iterative algorithms in non-convex regimes [22].
The following theorem shows convergence of the closed-loop system, combining Levels 1,2,3 in Figure 2.
Theorem 3.10 (Closed-loop scheme convergence).
Let Assumptions 1-5 hold. For $\eta\in(0,\frac{1}{2L^{\prime}})$, $\mu^{*}=2n^{1/4}\sqrt{{\epsilon_{g}}/{L_{j}}},j\in\{p,x\}$, the sequence $\{p^{k}\}_{k\in\mathbb{N}}$ generated by (16) satisfies
@@
where $K_{1}$ is given by
$K_{1}=\frac{6}{1-2\eta L^{\prime}}\Bigg{\{}\underset{\kappa_{1}}{\underbrace{\frac{\varphi(p^{0},h(p^{0},d))-\varphi^{*}}{3\eta|\mathcal{T}|}}}+\frac{\sigma_{\rm pe}^{2}}{\eta^{2}}+4\Bigg{[}\underset{\kappa_{2}}{\underbrace{\gamma^{2}L^{2}\epsilon_{x}^{2}}}+$
$\underset{\kappa_{3}}{\underbrace{n^{3/2}\epsilon_{g}\!(L_{p}\!\!+\!\!L^{2}L_{x})}}\!\Bigg{]}\!\!\!+\!\!{2(2n\gamma^{2}\!\!+\!\!M_{x}^{2}\!\!+\!\!4n^{3/2}\epsilon_{g}L_{x})}\frac{\sum_{\underset{l\leq k}{l\in\mathcal{T}}}\mathbb{E}[\lVert e^{l}\rVert^{2}]}{|\mathcal{T}|}\!\!\Bigg{\}},$
with $L^{\prime}=L_{p}+L^{2}(L_{x}+2)$ denoting the Lipschitz constant of the gradient $\Phi(\cdot,h(\cdot,d))$, and $\varphi^{*}~{}=~{}\underset{p\in[-1,1]^{n}}{\mathrm{inf}}\varphi(p,h(p,d))~{}>~{}-n$ the cost function value at a locally optimal solution.
@@
In Theorem 3.10, we use the expectation of the second moment on the fixed-point residual mapping as the convergence measure. This metric reflects the average performance of Algorithm 1. Moreover, we note that the term $\lVert\mathcal{G}(p^{k})\rVert$ approaches zero as $p^{k}$ approaches a locally optimal solution [43]. Note that an upper bound on the minimum of $\lVert\mathcal{G}(p^{k})\rVert$ can be inferred from (19), through the inequality $\mathbb{E}\Big{[}\underset{l\in\mathcal{T}}{\mathrm{min}}\lVert\mathcal{G}(p^{l})\rVert^{2}\Big{]}\leq\mathbb{E}\Big{[}\frac{1}{|\mathcal{T}|}\sum_{l\in\mathcal{T}}\lVert\mathcal{G}(p^{l})\rVert^{2}\Big{]}=\frac{1}{|\mathcal{T}|}\mathbb{E}\Big{[}\sum_{l\in\mathcal{T}}\lVert\mathcal{G}(p^{l})\rVert^{2}\Big{]}$, see [19, Remark 5]. Also, note that the theoretical bound in (19) is computed only over the trigger time instances. Since the positions between two consecutive trigger time instances do not change, the term $\lVert\mathcal{G}(p^{k})\rVert$ remains the same. Thus, in order to analyze convergence, it is sufficient to compute the fixed-point residual only at the trigger time instances.
The bound $K_{1}$ shows that achieving an accurate solution of (5), i.e., a local minima $(p^{*},h(p^{*},d))$, is limited by the deviation of the initial cost at $(p^{0},h(p^{0},d))$ from the optimal one $\varphi^{*}$ (i.e. term $\kappa_{1}$), the polarization and engagement gradient estimation errors (i.e. terms $\kappa_{2},\kappa_{3}$, respectively) including both opinion and clicking behaviour estimation errors, the variance of the dither signal $\sigma_{\rm pe}^{2}$, and the sensitivity estimation error variance $\mathbb{E}[\lVert e^{l}\rVert^{2}]$.
@@
As $|\mathcal{T}|\rightarrow\infty$, the term $K_{1}$ approaches $K_{2}$, given by
[…]
$4(\kappa_{2}+\kappa_{3})+\frac{\sigma_{\rm pe}^{2}}{\eta^{2}}\Bigg{\}}$
From $K_{2}$, we observe that the achievement of an asymptotically accurate solution of (5) at a local minima $(p^{*},h(p^{*},d))$ is limited by a combination of gradient estimation errors that already include opinion and clicking behaviour estimation (i.e. terms $\kappa_{2},\kappa_{3}$) and sensitivity estimation error, along with the variance of the dither signal $\sigma_{\rm pe}^{2}$. The term $\kappa_{1}$ approaches zero asymptotically. Moreover the upper-bound of the average of the variance of sensitivity estimation error over the trigger time instances approaches $C_{f}$ asymptotically.
4 Numerics

### v2

@@ 3 Recommender system design
We approach Problem 1 by designing a dynamic feedback controller inspired by the projected-gradient descent algorithm in [39, Section 3A]. The resulting recommender system dynamically generates recommendations as
@@
    Estimation of real-time opinions and users’ clicking behaviour via supervised learning;
@@ 3.1 Level I: Opinions & Clicking Behaviour Estimation
The opinions and clicking behaviours are estimated using an Artificial Neural Network (ANN).
We estimate the opinions by relying on implicit feedback from the users, i.e., the CTR $y\approx\mathbb{E}\big{[}\mathcal{B}(g(p,x))\big{]}$ of the users on recommendations with position $p$. Thus, we seek a map to extract the opinions from the clicks and the positions, $x=\beta(y,p)$. To bound the resulting opinions and clicking behaviour estimation error from the ANN, we work under the following regularity assumptions [21, 40].
@@
    There exists a continuous map $\beta:[0,1]^{n}\times[-1,1]^{n}\to[-1,1]^{n}$ is such that $\beta(y,p)=x+\theta(x)$ with $\lVert\theta(x)\rVert\leq\theta_{x}$, for all $x\in\mathcal{X}\subseteq[-1,1]^{n}$, and $\theta_{x}<\infty$; moreover, the image $\beta(\mathcal{Y},\mathcal{P})$ is compact for any $\mathcal{Y}\subseteq[0,1]^{n},\mathcal{P}\subseteq[-1,1]^{n}$.
@@
    There exists a constant $\alpha_{y}<\infty$ such that the composite function takes the form $g(p,\beta(y,p))=y+\nabla_{x}g(p,x)^{\top}\theta(x)+\alpha(y)$, with $\lVert\alpha(y)\rVert\leq\alpha_{y}$, for all $y\in\mathcal{Y}$. $\square$
Assumption 3 i) guarantees the map $\beta(y,p)$ to provide an estimate on users’ opinions upper bounded by $\theta_{x}$. Similarly, Assumption 3 ii), guarantees the (opinion) map $\beta(y,p)$ to (implicitly) provide an estimate on clicking behavior upper bounded by $M_{x}\theta_{x}+\alpha_{y}$, with $M_{x}$ resulting from Assumption 2. We will qualitatively describe how these upper bound errors can be tightened in Section 3.4.
@@ 3.1.1 Opinion Estimation
For each user $i\in[n]$, the input and output training sets are
$\begin{bmatrix}\mathcal{X}_{p}\\ \mathcal{X}_{y,i}\end{bmatrix}\in\begin{Bmatrix}[-1,1]^{n\times m}\\ [0,1]^{1\times m}\end{Bmatrix},\mathcal{X}_{x,i}\in[-1,1]^{1\times m},$
respectively. Given that users interact with each other, the recommendations to users connected to user $i$ will indirectly influence user $i$’s opinion $x_{i}$. However, $x_{i}$ is not directly affected by the CTR of other users. To model this, we design the input layer with $n+1$ neurons for each user: $n$ neurons associated with $p$ and one neuron associated with $y_{i}$. The output layer comprises a single neuron to represent $x_{i}$. Additionally, we incorporate an intermediate layer with $n+2$ neurons. The intermediate layer employs a hyperbolic tangent activation function $\tanh(x):\mathbb{R}\rightarrow[-1,1]$, while the output layer uses an identity activation function.
@@
Under Assumption 3, the opinion estimation error is upper bounded as $\lVert e_{x}\rVert\leq\epsilon_{x}$ with
@@
where $\underset{y\in\mathcal{X}_{y},p\in\mathcal{X}_{p}}{\mathrm{sup}}\lVert\beta(y,p)-\hat{\beta}(y,p)\rVert_{\infty}$ represents the maximum opinion estimation error during training of the ANN, $\omega_{\beta_{i}}(\gamma_{x})$ is the minimum modulus of continuity of $\beta_{i}$ on the training set as in Definition 1, and $v_{0,i}$ denotes the bias from the hidden layer to the output layer of the ANN on user $i$.
@@ 3.1.2 Clicking Behaviour Estimation
For each user $i\in[n]$, the input and output training sets are $\begin{bmatrix}\mathcal{X}_{p,i}\\ \mathcal{X}_{x,i}\end{bmatrix}\in[-1,1]^{2\times m}$, $\mathcal{X}_{y,i}\in[0,1]^{1\times m}$, respectively. We then construct an ANN similarly as for opinion estimation [^1] We omit the details due to space limitations and refer the interested reader to for a detail discussion.. {arxiv} To estimate the clicking behaviour, we proceed similarly as for opinion estimation. For each user $i\in[n]$, the input and output training sets are
$\begin{bmatrix}\mathcal{X}_{p,i}\\ \mathcal{X}_{x,i}\end{bmatrix}\in[-1,1]^{2\times m},\mathcal{X}_{y,i}\in[0,1]^{1\times m},$
respectively. Given the recommender system’s personalization of positions to the user’s interests, the $i$-th user’s CTR $y_{i}$ only depends on their steady-state opinion $x_{i}$ and the provided position $p_{i}$. Hence, we have two input neurons ($x_{i},p_{i}$) and one neuron in the output layer ($y_{i}$). Further, we use $3$ intermediate layers with $5$ neurons. The hyperbolic tangent and the identity are used as activation functions at the intermediate layers and the output layer, respectively.
@@
Under Assumptions 23, the clicking behaviour estimation error is upper-bounded as $\lVert e_{y}\rVert\leq\epsilon_{g}$, with
@@ 3.2 Level II: Online Sensitivity Learning
To estimate the sensitivity $\nabla_{p}h(p,d)$ in (7), we develop a Kalman filter, similar to [19]. Given that opinion dynamics and recommendations evolve over time scales that are comparable in magnitude, in order to make our OFO scheme [41] applicable, we introduce an auxiliary mechanism that guarantees time-scale separation between the opinion dynamics and the recommender system updates, see Remark 1.
@@ 3.2.1 Kalman Filter for Sensitivity Learning
The measurement model (11) is obtained from a first-order Taylor approximation of the steady-state map around $p^{k-1}$, i.e., $h(p^{k},d)\approx h(p^{k-1},d)+\nabla_{p}h(p^{k-1},d)^{\top}(p^{k}-p^{k-1})$. With this intuition, it can be seen that the measurement model is a linear time-variant approximation of the temporal evolution of the sensitivity $\nabla_{p}h(p^{k},d)$. Note that, the steady-state $h(p^{k},d)$ is reached only asymptotically. Thus, it is common practice [19] to use instead $\Delta x^{k+1,k}_{ss}=x^{k+1}-x^{k}$ as the measurement in (11). $\square$
@@
The trigger mechanism is introduced for the recommender system to update positions close to steady state, see Remark 1, and to collect a sufficient number of clicks to ensure a proper estimate of the CTR needed for opinion estimation. The trigger mechanism is turned on at integer multiples $i\in\mathbb{N}$ of the time period $T$. We define $\mathcal{T}$ as the set of all trigger time instances, with $\tau_{i}\in\mathcal{T}$ representing the most recent trigger time instance prior to $k$. Note that the update (3.2.1) uses the opinion estimate $\hat{x}$ introduced in §3.1.1 rather than the real users’ opinion, with $\hat{x}^{k+1}=\hat{\beta}(y^{k},p^{k})$, where $y^{k}:=\sum_{t=\tau_{i}}^{k}c^{t}/(k-\tau_{i}+1)$ is the CTR since the most recent trigger time instant. Finally, the Kalman filter gain $K^{k}$ in (3.2.1) is given by
@@
The assumption of white noise for the process model is standard in the context of sensitivity learning in feedback optimization [19]. Intuitively, the process perturbation determines the degree of trust one puts on the sensitivity estimates. It is also reasonable to assume that the external influence is uncorrelated among users in certain cases, for example the extended FJ model in (2) where $d=x_{0}$. We also note that since $\hat{x}$ is estimated using the ANN, the steady-state estimation error $e_{x}$ is not correlated with the process or measurement noises in (10), (11). Under Assumption 4, the covariances simplify as $Q^{k}=(\sigma_{q}^{k})^{2}I_{n^{2}}$ and $R^{k}=(\sigma_{r}^{k})^{2}I_{n}$, for some $\sigma_{q}$, $\sigma_{r}>0$. The tuning of $\sigma_{q},\sigma_{r}$, is based on heuristics and is described in Appendix 5.4.
@@
In practice, this assumption can be satisfied by introducing a dither signal in the update rule (6).
@@ 3.3.1 Gradient Estimation
where $\nabla_{j}\hat{\varphi}^{\rm ctr}_{i},j\in\{p,x\}$ denotes the $i$-th entry of the gradient, $e_{i}\in\mathbb{R}^{n}$ refers to the $i$-th vector of the canonical basis of $\mathbb{R}^{n}$, and $\mu$ is a smoothing parameter. Note that in the two-point gradient estimation, we do not actuate the opinion dynamics twice but rather use the engagement cost function estimate $\hat{\varphi}^{\rm ctr}(p,x)=-1_{n}^{\top}\hat{g}(p,x)$ as described in §3.1.2.
@@ 3.4 Closed-loop Convergence Guarantees
Here, we establish convergence guarantees for the sensitivity estimation process (3.2.1) and for the closed-loop interconnection between the opinions and recommendations. Before stating the formal results, we make some important observations regarding our opinion estimation technique and its implication on the closed-loop convergence analysis.
In our problem setup, the measurement information at each time instant (i.e., a single click/no click data), is not rich enough to produce accurate estimates for the users’ real-time opinion. As shown in Algorithm 1, the opinion estimation is instead carried out using the CTR recorded on multiple recommendations with a constant position since the previous trigger instant. In light of the exponential stability of the opinion dynamics (i.e., Assumption 1), the resulting estimate at the end of each trigger period is, in practice, close to an estimate of the steady-state opinion $h(p^{k},d)$. This approach is therefore substantially different from real-time state estimation [21]. Thus, as we directly estimate steady-state opinions at each trigger instant, we can greatly simplify our closed-loop stability analysis by treating the opinion dynamics as a static mapping. Also note that as the trigger period $T$ increases, the resulting CTR offers a better approximation of the CTR in expectation $\mathbb{E}\big{[}g(p,x)\big{]}$. The modeling errors $\theta_{x},\alpha_{y}$ in Lemma 23.2 exist due to a finite time horizon $T$ and reduces as $T$ increases.
@@
To quantify performance on the recommendations, we use the fixed-point residual mapping [43, Eq. (5)]
$\mathcal{G}(p):=\frac{1}{\eta}\Big{(}p-\Pi_{[-1,1]^{n}}\big{[}p-\eta\Phi(p,h(p,d))\big{]}\Big{)}.$ (18)
The fixed-point residual mapping is zero at any critical point of (5), and is a common metric to quantify convergence of iterative algorithms in non-convex regimes [26].
The following theorem shows the convergence of the closed-loop system, combining Levels I, II, and III in Figure 2, using the second moment of the fixed-point residual mapping. Since the problem is non-convex, the trajectories of $\{p^{k}\}_{k=0}^{N}$ vary across each trial. Thus, it is essential to evaluate $\lVert\mathcal{G}(p^{k})\rVert$ with its expectation. Moreover, we compute the average of this metric over all the trigger time instances to reflect the average performance of Algorithm 1 over the time in which recommendations are provided. The theoretical bound is only computed over the trigger time instances since the positions remain constant between two trigger instances, and so does the residual $\lVert\mathcal{G}(p^{k})\rVert$.
Theorem 3.10 (Closed-loop convergence).
Let Assumptions 15 hold. For any $\eta\in(0,\frac{1}{2L^{\prime}})$ and $\mu^{*}=2n^{1/4}\sqrt{{\epsilon_{g}}/{L_{j}}},j\in\{p,x\}$, the sequence $\{p^{k}\}_{k\in\mathbb{N}}$ generated by (16) satisfies
@@
where the upper bound $K_{1}$ is given by
$K_{1}=\frac{6}{1-2\eta L^{\prime}}\Bigg{(}\underset{\kappa_{1}}{\underbrace{\frac{\varphi(p^{0},h(p^{0},d))-\varphi^{*}}{3\eta|\mathcal{T}|}}}+\frac{\sigma_{\rm pe}^{2}}{\eta^{2}}+4\Bigg{(}\underset{\kappa_{2}}{\underbrace{\gamma^{2}L^{2}\epsilon_{x}^{2}}}+$
$\underset{\kappa_{3}}{\underbrace{n^{3/2}\epsilon_{g}\!(L_{p}\!\!+\!\!L^{2}L_{x})}}\!\Bigg{)}\!\!\!+\!\!{2(2n\gamma^{2}\!\!+\!\!M_{x}^{2}\!\!+\!\!4n^{3/2}\epsilon_{g}L_{x})}\frac{\sum_{\underset{l\leq k}{l\in\mathcal{T}}}\mathbb{E}[\lVert e^{l}\rVert^{2}]}{|\mathcal{T}|}\ \!\!\Bigg{)},$
In there, $L^{\prime}=L_{p}+L^{2}(L_{x}+2)$ is the Lipschitz constant of the gradient $\Phi(\cdot,h(\cdot,d))$, and $\varphi^{*}={\mathrm{inf}}_{p\in[-1,1]^{n}}\varphi(p,h(p,d))>-n$ is the cost function value at a locally optimal solution.
@@
Note that an upper bound on the minimum of $\lVert\mathcal{G}(p^{k})\rVert$ can be inferred from (19), through the inequality $\mathbb{E}\Big{[}\underset{l\in\mathcal{T}}{\mathrm{min}}\lVert\mathcal{G}(p^{l})\rVert^{2}\Big{]}\leq\mathbb{E}\Big{[}\frac{1}{|\mathcal{T}|}\sum_{l\in\mathcal{T}}\lVert\mathcal{G}(p^{l})\rVert^{2}\Big{]}=\frac{1}{|\mathcal{T}|}\mathbb{E}\Big{[}\sum_{l\in\mathcal{T}}\lVert\mathcal{G}(p^{l})\rVert^{2}\Big{]}$, see [23, Remark 5]. Thus, intuitively, the bound in (19) describes the maximum possible minimum gap between the trajectory $\{p^{k},h(p^{k},d)\}_{k=0}^{N}$ obtained using Algorithm 1 and a local minima $(p^{*},h(p^{*},d))$.
The bound $K_{1}$ shows that achieving an accurate solution of (5), i.e., a local minima $(p^{*},h(p^{*},d))$, is limited by the deviation of the initial cost at $(p^{0},h(p^{0},d))$ from the optimal one $\varphi^{*}$ (i.e., the term $\kappa_{1}$), the polarization and engagement gradient estimation errors (i.e., the terms $\kappa_{2},\kappa_{3}$, respectively) including both opinion and clicking behaviour estimation errors, the variance of the dither signal $\sigma_{\rm pe}^{2}$, and the sensitivity estimation error variance $\mathbb{E}[\lVert e^{l}\rVert^{2}]$.
@@
As $|\mathcal{T}|\rightarrow\infty$, the upper bound $K_{1}$ in in Theorem 3.10 approaches $K_{2}$, given by
$K_{2}=\frac{6}{1-2\eta L^{\prime}}\Bigg{(}$ $2C_{f}(2n\gamma^{2}+M_{x}^{2}+4n^{3/2}\epsilon_{g}L_{x})+$
$4(\kappa_{2}+\kappa_{3})+\frac{\sigma_{\rm pe}^{2}}{\eta^{2}}\Bigg{)}$
Proof 3.13.
Refer to Appendix 5.15.
From $K_{2}$, we observe that the asymptotic accuracy is limited by a combination of gradient estimation errors, which include opinion and clicking behaviour estimation (i.e., the terms $\kappa_{2},\kappa_{3}$), the sensitivity estimation error, and the variance of the dither signal $\sigma_{\rm pe}^{2}$. Moreover, the upper-bound of the average of the variance of sensitivity estimation error over the trigger time instances approaches $C_{f}$, the asymptotic sensitivity error variance.
4 Numerical Case Studies